As a person inhales, air moves down the windpipe (bronchus), through a constriction where the air speed doubles. If the air is travelling 13 cm/s before the constriction and we treat air as an incompressible fluid,

determine the pressure drop in the constriction. Use the density of air as 1.29 kg/m3.
Pa
Answer #1

From the question the initial speed of the air is v1 = 13 cm/s = 0.13 m/s
final speed of the air flow is v2 = 2*v1 = 2*0.13 = 0.26 m/s
density of air is ρ = 1.29 kg/m ^3
Using Bernouli's theoremP1+(1/2)ρ(v1)2 =P2+(1/2)ρ(v2)2
So pressure drop P1-P2 =(1/2)ρ(v2)2 -(1/2)ρ(v1)2   -----1
putting the values in equation 1 and solve for P1 -P2 .
   P1 -P2 = (1/2)*1.29*[(0.26)^2 -( 0.13 )^2]
             = 0.033 Pa